OpenStudy (anonymous):

Use the quadratic formula to solve 16x2 - 32x + 16 = 0

OpenStudy (anonymous):

the quadratic formula is \[\huge x\; =\; \frac{-b\; \pm \;\sqrt{\;b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

Let us get to the facts The general equation of a quadrtic is ax^2+bx+c=0 Drawing analogies with the above equation a=16, b=-32 and c=16 Using \[x=-b \pm \sqrt{b^2-4ac}\]

OpenStudy (anonymous):

you will get x = 16